For 2017, country ABC has the following data (all figures are in billion), AE = $1720, govt. expenditure = $500, Consumption expenditure = $1000, Investment expenditure = $200, Exports = $75. Calculate imports by this country in 2017. 


1
Expert's answer
2021-03-17T08:33:07-0400

"AE = C + I + G + NX"

Where, AE = Aggregate expenditure

           C = Consumption expenditure

           I = Investments expenditure

           G = Government expenditure

           NX = Net exports (Exports - Imports)

Thus,

1720 = 1000 + 200 + 500 + 75 – M

1720 = 1775 – M

Imports, M = 1775 – 1720 = 55